OpenStudy (anonymous):

Complete each conjecture. A rule in terms of n for the sum of the first n EVEN positive integers is __?__

OpenStudy (dan815):

2=1^2+1 2+4=6=2^2+2 2+4+6=12=3^2+3 .................. n^2+n

OpenStudy (e.mccormick):

OpenStudy is for tutoring and learning. So, lets say you make a post, someone comes to tutor you, and you just try to get them to give you the answer... close it and re-post because they did not direct answer... etc. That is not a proper use of OpenStudy. Instead, you need to work with the people so they can lead you towards the right path. Then you can take that path and find the answer for yourself. Once you have learned the concept, you should not need to post the same or similar questions over and over, which is what has caused dan some frustrations in working with you.